# Dexterous Manipulation through Imitation Learning: A Survey

### v1

@@ III-E1 Description
Continual Imitation Learning (CIL) is an approach that integrates continual learning with imitation learning, aiming to enable agents to continuously acquire and adapt skills by imitating expert behaviors in dynamically changing environments. Specifically, in the initial phase, the agent learns fundamental skills from expert demonstrations. In subsequent phases, the agent incrementally accumulates knowledge, adapts to new tasks or environments, and mitigates the risk of forgetting previously acquired skills.
@@ IV-A Two-fingered Traditional Gripper
Two-fingered grippers are widely used for their reliability, simplicity, and ease of control. Typically driven by a single actuator with one degree of freedom (DoF), they are cost-effective and suitable for repetitive tasks requiring consistency [103]. For example, [104] demonstrated a Franka robot with such a gripper performing tasks like setting a breakfast table. Similarly, Kim et al. [105] used a two-fingered gripper for behavior cloning with gaze prediction, and a tendon-driven variant in [106] showed the ability to grasp diverse household objects.
Recent works have further extended gripper capabilities through large-scale imitation datasets such as MIME [107], RH20T [108], Bridge Data [109], and Droid [110]. Dual-arm systems also enhance manipulation by coordinating two grippers. For instance, [111] achieved banana peeling through dual-action imitation learning. More complex tasks such as shrimp cooking, cloth folding, and dishwashing were demonstrated in Mobile ALOHA [112] and UMI [113].
@@ IV-B Multi-fingered Anthropomorphic Hand
Figure 3: Examples of multi-fingered anthropomorphic hands: (a) Shadow Dexterous Hand[119]; (b) Awiwi Hand [120]; (c) Biomimetic Hand [121]; (d) ILDA Hand [117]; (e) Hu et al.’s robotic hand [118]; (f) INSPIRE-ROBOTS RH56 Dexterous Hand[122]; (g) OYMotion OHand[123]; (h) PUT-Hand [124]; (i) Allegro Hand[125]; (j) Faive Hand [126]; (k) Tesla Optimus Hand[127]; (l) Utah/M.I.T. Dexterous Hand [128].
@@ IV-B1 Tendon-driven Approach
Tendon-driven hands use cable transmissions to actuate joints, mimicking human tendons. This design allows compact structure, multiple DoFs, and high dexterity, making it a common choice in anthropomorphic hand development. To accommodate high DoFs, actuators are often remotely located in the forearm.
@@ IV-C Three-fingered Robotic Claw: A Trade-off Solution
As a compromise between the minimalistic two-fingered gripper and complex multi-fingered anthropomorphic hands, the three-fingered robotic claw offers a functional middle ground. Though not anatomically human-like, three fingers are sufficient for executing common grasp types such as cylindrical and spherical power grasps [199], and can support a subset of in-hand manipulation tasks.
Numerous three-fingered claws have demonstrated impressive capabilities. For example, Shadow’s DEX-EE [204] (see Fig. 4(a)) and the TRX Hand [207] exhibit high robustness and dexterity. The BarrettHand [208] (see Fig. 4(b)) achieves adaptive grasping through underactuation. Tendon-driven designs like the i-HY Hand [199] (see Fig. 4(c)) and Model O [209] enable actions such as pivoting and precision transitions. Systems such as DClaw [210] and TriFinger [211] are capable of performing fine manipulation tasks via reinforcement learning. Other novel architectures include linkage-based [212], motor-multiplexed [213], and link-belt-integrated claws [214], each offering different characteristics.
@@ V Teleoperation Systems and Data Collection
Teleoperation systems provide a robust interface for human-robot collaboration, benefiting from directly making robot behaviors comply with human-level intelligence, which refers to human-in-the-loop. This approach is highly intuitive since humans’ extensive knowledge and experience empower them to make informed judgments on diverse tasks across complex scenes and to promptly adjust strategies in response to feedback. Due to this usability, teleoperation is widely applied in various fields. Additionally, by collecting data on the robot’s states and corresponding actions during teleoperation, datasets can be constructed to perform end-to-end imitation learning.
@@ V-A Teleoperation Systems for Dexterous Manipulation
To accurately convey human operators’ intentions to robotic systems, previous works have employed a wide range of human-robot interaction devices. Human operators with work experience can easily identify the current state of a robot through the image, however, accurately translate human instructions to robot actions remains a challenge. Some traditional controllers act on this: 1) joysticks [218] 2) haptic devices [219]; However, manipulation tasks often involve delicate movements and complex interactions, such as grasping, moving, and positioning small or irregularly shaped objects. These tasks necessitate devices that can offer dexterous interfaces to ensure the safety and efficacy of the robot’s actions. Precision and real-time feedback are crucial. Commonly used devices include: 1) cameras [18, 220, 17, 19, 221, 222, 223, 54, 20]; 2) mocap gloves [224, 225, 226, 227, 228, 229]; 3) VR/AR controllers [230, 231, 14, 232, 233, 54, 234, 27, 235, 236, 237, 238]; 4) exoskeletons and bilateral systems [239, 240, 241, 53, 242, 243].
@@ V-A1 Vision-based Teleoperation Systems
Recently, advancements in computer vision have led to the development of vision-based teleoperation systems. However, their accuracy in capturing hand movements is often compromised by factors such as occlusion, lighting, resolution, background, and inaccurate 3D estimation issues. Several methods have been proposed for robust hand pose estimation and reliable mapping to the robot end-effector. Li et al. [222] developed a vision-based teleoperation system by training TeachNet on pairs of images of human hands and simulated robots to form mappings between a human hand and a robotic Shadow Hand in the latent space. Dexpilot [18] utilized a calibrated multi-camera system to estimate hand poses to teleoperate an Allegro Hand. Riemannian Motion Policies (RMPs) are employed to compute the Cartesian pose of the hand, facilitating hand-arm motion control. Subsequent approaches, such as Robotic Telekinesis [17] and DIME [54], which simplified requirements to a single RGB camera, thus reducing the need for calibration. This is achieved through a general mapping method between humans and robots that have different kinematic structures. Additionally, Robotic telekinesis [17] adjusts the position and orientation of the end-effector relative to its base using the relative position and direction of the human wrist to the torso, enabling the teleoperate both arm and hand. However, these methods still suffer from occlusion issues due to the single fixed camera setting. To solve this problem, Transteleop [223] introduced a system that utilizes real-time active vision with a depth camera mounted on the end-effector of the remote UR5 robot arm. During teleoperation, this robot arm can reposition the camera to enhance its field of view and improve hand pose estimation accuracy.
@@ V-A2 Mocap Gloves
Motion capture systems typically utilize stable hardware devices such as multi-camera setups with markers, IMU sensors, and RGB-D cameras. These devices are robust against changes in lighting, occlusion, and complex backgrounds. Mocap gloves collect human hand motion data directly via sensors, enabling ideal real-time performance and significantly improving data collection efficiency in teleoperation. Although motion capture gloves are expensive, they provide precise hand tracking [226]. Wang et al. [16] introduced DexCap, a portable motion capture system. It includes a mocap glove for accurate finger joint tracking, a single-view camera for 6-DoF wrist pose tracking, and an RGB-D LiDAR camera for observing the surrounding 3D environment. With this precise 3D hand motion data, the proposed DexIL system can effectively learn bimanual dexterous manipulation skills. The remote system features two Franka Emika robotic arms, each outfitted with a LEAP dexterous robotic hand. Similarly, Mosbach et al. [245] used the SenseGlove DK1, a force-feedback glove, to capture hand joint movements, with hand tracking facilitated by a camera mounted on a headset.
@@ V-A4 Exoskeleton and Bilateral Systems
The majority of aforementioned methods focus on manipulating the robot’s end-effector in task space in Cartesian coordinates. While setting the robot’s end-effector position is convenient, it has drawbacks. For robots with multiple degrees-of-freedom (multi-DoF), computationally demanding inverse kinematics (IK) calculations are required, which can be problematic in real-time control scenarios. These complexities may cause response delays and compromise operational precision. Furthermore, singularities in the motion trajectory may lead to indeterminate or nonexistent IK solutions, resulting in control failures. In the following sections, we examine studies that aim to synchronize human and robot movements in the joint space.
@@
Exoskeletons are wearable devices that gather and analyze user motion data. Fabian [239] developed a lightweight exoskeleton, DE VITO for measuring human arm movements to teleoperate the mobile robot DE NIRO[249]. Meanwhile, AirExo [240] presents a framework for whole-arm dexterous manipulation adaptable to different robot arms using interchangeable 3D-printed components for robots divergent in morphology.
@@
Another approach involves a bilateral framework, where movements of the leader robot are mirrored by the follower robot. Any resistance or force encountered by the follower is communicated back to the leader, enabling precision and tactile sensation tasks. Kim et al. [241] develops a controller with Denavit-Hartenberg (DH) parameters matching the teleoperated dual-arm robot, alongside a calibration method to reduce gravitational errors. For demonstrations without real robots, the controller uses force/torque (F/T) sensors identical to those in real robots to provide force feedback. Recently, ALOHA [53] utilizes structurally analogous robotic arms with identical joint spaces for teleoperation, employing cost-effective ViperX [250] aarms as follower robots and WidowX [251] of comparable size serve as leaders to enhance control capabilities
Expanding upon this concept, Mobile ALOHA [242] integrates the system with an Automated Guided Vehicle (AGV) to establish a whole-body teleoperation system. GELLO [243] has further reduced costs by replacing the real robotic arms at the local site with scaled kinematically equivalent 3D-printed parts and off-the-shelf motors, achieving one-to-one joint mapping.
@@
Differing from specific robot methods, AnyTeleop [19] introduced a unified system supporting multiple robot arms and dexterous hands through a general motion retargeting method, which maps robot hands to human fingers. This system supports different arms by generating trajectories based on the estimated Cartesian end-effector pose. ACE [221] developed a cross-platform visual-exoskeleton teleoperation system compatible with diverse robot hardware, including various end-effectors such as grippers and multi-finger hands, offering flexibility. Its exoskeleton arm features high-resolution encoders for precise joint position readings, ensuring accurate end-effector tracking.
@@ V-B1 Datasets
MIME [252]is a large-scale dataset contains 8,260 human-robot demonstrations across 20 diverse tasks ranging from simple tasks like pouring to difficult tasks like stacking objects. It includes both videos of human demonstrations and kinesthetic trajectories for robots.
RH20T [253] dataset encompasses over 110,000 multimodal robotic manipulation sequences collected using intuitive teleoperation interfaces equipped with force-torque sensors and haptic feedback. It captures visual, tactile, audio, and proprioceptive data alongside human demonstration videos, facilitating one-shot imitation learning across diverse tasks and robotic configurations. This dataset aims to advance robotic skill acquisition in unstructured environments through enhanced task and motion planning.
@@
DRIOD [256] surpasses other datasets collected purely via human teleoperation with its unparalleled scene diversity and task variety. It is a diverse robot imitation learning dataset with 76,000 demonstration trajectories across 86 tasks and 564 scenes.
To address the problem that collecting extensive human demonstrations for imitation learning is always time-consuming and laborious. Several datasets apply data augmentation to demonstrations collected via human teleoperation. To illustrate, data augmentation is a technique that has been widely applied in various fields, particularly in computer vision and robotics. It is used to artificially expand the size and diversity of a training dataset by applying a series of transformations or modifications to the original data, which can improve the robustness and generalization capabilities of machine learning models trained on the augmented dataset.
@@
Several other datasets leverage the demonstration generation system to enhance imitation learning by expanding the dataset from a limited number of human demonstrations. In MimicGen [259], over 50,000 demonstrations are created across 18 tasks from approximately 200 human demonstrations, it synthesizes full demonstrations of diverse scene configurations, object instances, and robot arms by adapting object-centric manipulation behaviors to new contexts through trajectory transformations based on known object poses, enabling the training for complex, long-horizon, and high-precision tasks.
@@
Additionally, DexGraspNet [263] featuring 1.32 million grasps of 5,355 objects by ShadowHand fills the blank of large-scale, diverse, and high-quality datasets for dexterous grasping. It includes over 200 diverse grasps per object, validated for physical stability in simulation, enabling more effective imitation learning and benchmarking of robotic manipulation algorithms to achieve human-like dexterity and grasping capabilities.
@@ VI Challenges and Future Directions in Imitation Learning-Based Dexterous Manipulation
Imitation learning-based dexterous manipulation poses unique challenges due to the inherent complexities of both imitation learning and dexterous control. Despite significant advancements over the past decade, several challenges hinder its human-level dexterity and real-world applicability. This section discusses these challenges from multiple perspectives and explores future research directions.
@@ VI-A Data Collection and Generation
Data collection and generation for imitation learning-based dexterous manipulation poses several challenges, including heterogeneous data fusion, data diversity, high-dimensional data sparsity, and data collection costs:

### v2

@@ III-E1 Description
Continual Imitation Learning (CIL) is an approach that integrates continual learning with IL, aiming to enable agents to continuously acquire and adapt skills by imitating expert behaviors in dynamically changing environments. Specifically, in the initial phase, the agent learns fundamental skills from expert demonstrations. In subsequent phases, the agent incrementally accumulates knowledge, adapts to new tasks or environments, and mitigates the risk of forgetting previously acquired skills.
@@ IV-A Two-fingered Traditional Gripper
Two-fingered grippers are widely used for their reliability, simplicity, and ease of control. Typically driven by a single actuator with one DoF, they are cost-effective and suitable for repetitive tasks requiring consistency [103]. For example, [104] demonstrated a Franka robot with such a gripper performing tasks like setting a breakfast table. Similarly, Kim et al. [105] used a two-fingered gripper for behavior cloning with gaze prediction, and a tendon-driven variant in [106] showed the ability to grasp diverse household objects.
Recent works have further extended gripper capabilities through large-scale imitation datasets such as MIME [107], RH20T [108], Bridge Data [109], and Droid [110]. Dual-arm systems also enhance manipulation by coordinating two grippers. For instance, [111] achieved banana peeling through dual-action IL. More complex tasks such as shrimp cooking, cloth folding, and dishwashing were demonstrated in Mobile ALOHA [112] and UMI [113].
@@ IV-B Multi-fingered Anthropomorphic Hand
Figure 3: Examples of multi-fingered anthropomorphic hands: (a) Shadow Dexterous Hand[119]; (b) Awiwi Hand [120]; (c) Biomimetic Hand [121]; (d) ILDA Hand [117]; (e) Hu et al.’s robotic hand [118]; (f) INSPIRE-ROBOTS RH56 Dexterous Hand[122]; (g) OYMotion OHand[123]; (h) PUT-Hand [124]; (i) Allegro Hand[125]; (j) Faive Hand [126]; (k) Tesla Optimus Hand[127]; (l) Utah/MIT Dexterous Hand [128].
@@ IV-B1 Tendon-driven Approach
Tendon-driven hands use cable transmissions to actuate joints, mimicking human tendons. This design allows for compact structure, multiple DoFs, and high dexterity, making it a common choice in anthropomorphic hand development. To accommodate high DoFs, actuators are often remotely located in the forearm.
@@ IV-C Three-fingered Robotic Claw: A Trade-off Solution
As a compromise between the minimalist two-fingered gripper and complex multi-fingered anthropomorphic hands, the three-fingered robotic claw offers a functional middle ground. Though not anatomically human-like, three fingers are sufficient for executing common grasp types such as cylindrical and spherical power grasps [199], and can support a subset of in-hand manipulation tasks.
Numerous three-fingered claws have demonstrated impressive capabilities. For example, Shadow’s DEX-EE [204] (see Fig. 4(a)) and the TRX Hand [207] exhibit high robustness and dexterity. The BarrettHand [208] (see Fig. 4(b)) achieves adaptive grasping through underactuation. Tendon-driven designs like the i-HY Hand [199] (see Fig. 4(c)) and Model O [209] enable actions such as pivoting and precision transitions. Systems such as DClaw [210] and TriFinger [211] are capable of performing fine manipulation tasks via RL. Other novel architectures include linkage-based [212], motor-multiplexed [213], and link-belt-integrated claws [214], each offering different characteristics.
@@ V Teleoperation Systems and Data Collection
Teleoperation systems provide a robust interface for human-robot collaboration, benefiting from directly making robot behaviors comply with human-level intelligence, which refers to human-in-the-loop. This approach is highly intuitive since humans’ extensive knowledge and experience empower them to make informed judgments on diverse tasks across complex scenes and to promptly adjust strategies in response to feedback. Due to this usability, teleoperation is widely applied in various fields. Additionally, by collecting data on the robot’s states and corresponding actions during teleoperation, datasets can be constructed to perform end-to-end IL.
@@ V-A Teleoperation Systems for Dexterous Manipulation
To accurately convey human operators’ intentions to robotic systems, previous works have employed a wide range of human-robot interaction devices. Human operators with work experience can easily identify the current state of a robot through the image; however, accurately translating human instructions to robot actions remains a challenge. Some traditional controllers act on this: 1) joysticks [218] 2) haptic devices [219]; However, manipulation tasks often involve delicate movements and complex interactions, such as grasping, moving, and positioning small or irregularly shaped objects. These tasks necessitate devices that can offer dexterous interfaces to ensure the safety and efficacy of the robot’s actions. Precision and real-time feedback are crucial. Commonly used devices include: 1) cameras [18, 220, 17, 19, 221, 222, 223, 54, 20]; 2) mocap gloves [224, 225, 226, 227, 228, 229]; 3) VR/AR controllers [230, 231, 14, 232, 233, 54, 234, 27, 235, 236, 237, 238]; 4) exoskeletons and bilateral systems [239, 240, 241, 53, 242, 243].
@@ V-A1 Vision-based Teleoperation Systems
Recently, advancements in computer vision have led to the development of vision-based teleoperation systems. However, their accuracy in capturing hand movements is often compromised by factors such as occlusion, lighting, resolution, background, and inaccurate 3D estimation issues. Several methods have been proposed for robust hand pose estimation and reliable mapping to the robot end-effector. Li et al. [222] developed a vision-based teleoperation system by training TeachNet on pairs of images of human hands and simulated robots to form mappings between a human hand and a robotic Shadow Hand in the latent space. Dexpilot [18] utilized a calibrated multi-camera system to estimate hand poses to teleoperate an Allegro Hand. Riemannian Motion Policies (RMPs) are employed to compute the Cartesian pose of the hand, facilitating hand-arm motion control. Subsequent approaches, such as Robotic Telekinesis [17] and DIME [54], simplified requirements to a single RGB camera, thus reducing the need for calibration. This is achieved through a general mapping method between humans and robots that have different kinematic structures. Additionally, Robotic telekinesis [17] adjusts the position and orientation of the end-effector relative to its base using the relative position and direction of the human wrist to the torso, enabling the teleoperate both arm and hand. However, these methods still suffer from occlusion issues due to the single fixed camera setting. To solve this problem, Transteleop [223] introduced a system that utilizes real-time active vision with a depth camera mounted on the end-effector of the remote UR5 robot arm. During teleoperation, this robot arm can reposition the camera to enhance its field of view and improve hand pose estimation accuracy.
@@ V-A2 Mocap Gloves
Motion capture systems typically utilize stable hardware devices such as multi-camera setups with markers, IMU sensors, and RGB-D cameras. These devices are robust against changes in lighting, occlusion, and complex backgrounds. Mocap gloves collect human hand motion data directly via sensors, enabling ideal real-time performance and significantly improving data collection efficiency in teleoperation. Although motion capture gloves are expensive, they provide precise hand tracking [226]. Wang et al. [16] introduced DexCap, a portable motion capture system. It includes a mocap glove for accurate finger joint tracking, a single-view camera for 6-DoF wrist pose tracking, and an RGB-D LiDAR camera for observing the surrounding 3D environment. With this precise 3D hand motion data, the proposed DexIL system can effectively learn bimanual dexterous manipulation skills. The remote system features two Franka Emika robotic arms, each with a LEAP dexterous robotic hand. Similarly, Mosbach et al. [245] used the SenseGlove DK1, a force-feedback glove, to capture hand joint movements, with hand tracking facilitated by a camera mounted on a headset.
@@ V-A4 Exoskeleton and Bilateral Systems
The majority of the aforementioned methods focus on manipulating the robot’s end-effector in task space in Cartesian coordinates. While setting the robot’s end-effector position is convenient, it has drawbacks. For robots with multiple DoF, computationally demanding inverse kinematics (IK) calculations are required, which can be problematic in real-time control scenarios. These complexities may cause response delays and compromise operational precision. Furthermore, singularities in the motion trajectory may lead to indeterminate or nonexistent IK solutions, resulting in control failures. In the following sections, we discuss studies that aim to synchronize human and robot movements in the joint space.
@@
Exoskeletons are wearable devices that gather and analyze user motion data. Fabian [239] developed a lightweight exoskeleton, DE VITO, for measuring human arm movements to teleoperate the mobile robot DE NIRO[249]. Meanwhile, AirExo [240] presents a framework for whole-arm dexterous manipulation adaptable to different robot arms using interchangeable 3D-printed components for robots divergent in morphology.
@@
Another approach involves a bilateral framework, where the movements of the leader robot are mirrored by the follower robot. Any resistance or force encountered by the follower is communicated back to the leader, enabling precision and tactile sensation tasks. Kim et al. [241] develop a controller with Denavit-Hartenberg (DH) parameters matching the teleoperated dual-arm robot alongside a calibration method to reduce gravitational errors. For demonstrations without real robots, the controller uses force/torque (F/T) sensors identical to those in real robots to provide force feedback. Recently, employing cost-effective arms with comparable size as leaders and followers, ALOHA [53] utilizes structurally analogous robotic arms with identical joint spaces for teleoperation. Expanding upon this concept, Mobile ALOHA [242] integrates the system with an Automated Guided Vehicle (AGV) to establish a whole-body teleoperation system. GELLO [243] reduces costs by replacing the real robotic arms at the local site with scaled kinematically equivalent 3D-printed parts, achieving one-to-one joint mapping. After that, UMI [113] further eliminates the need for physical robot arms by using hand-held grippers, providing a more portable interface for in-the-wild data collection.
@@
Differing from specific robot methods, AnyTeleop [19] introduced a unified system supporting multiple robot arms and dexterous hands through a general human-robot hand retargeting method. This system supports different arms by generating trajectories based on the estimated Cartesian end-effector pose. ACE [221] developed a cross-platform visual-exoskeleton teleoperation system compatible with diverse robot hardware, including various end-effectors such as grippers and multi-finger hands, offering flexibility. Its exoskeleton arm features high-resolution encoders for precise joint position readings, ensuring accurate end-effector tracking.
@@ V-B1 Datasets
MIME [250]is a large-scale dataset containing 8,260 human-robot demonstrations across 20 diverse tasks ranging from simple tasks like pouring to difficult tasks like stacking objects. It includes both videos of human demonstrations and kinesthetic trajectories for robots.
RH20T [251] dataset encompasses over 110,000 multimodal robotic manipulation sequences collected using intuitive teleoperation interfaces equipped with force-torque sensors and haptic feedback. It captures visual, tactile, audio, and proprioceptive data alongside human demonstration videos, facilitating one-shot IL across diverse tasks and robotic configurations. This dataset aims to advance robotic skill acquisition in unstructured environments through enhanced task and motion planning.
@@
DROID [254] surpasses other datasets collected purely via human teleoperation with its unparalleled scene diversity and task variety. It is a diverse robot imitation learning dataset with 76,000 demonstration trajectories across 86 tasks and 564 scenes.
To address the problem that collecting extensive human demonstrations for IL is always time-consuming and laborious, several datasets apply data augmentation to demonstrations collected via human teleoperation. To illustrate, data augmentation is a technique that has been widely applied in various fields, particularly in computer vision and robotics. It is used to artificially expand the size and diversity of a training dataset by applying a series of transformations or modifications to the original data, which can improve the robustness and generalization capabilities of machine learning models trained on the augmented dataset.
@@
Several other datasets leverage the demonstration generation system to enhance IL by expanding the dataset from a limited number of human demonstrations. In MimicGen [257], over 50,000 demonstrations are created across 18 tasks from approximately 200 human demonstrations; it synthesizes full demonstrations of diverse scene configurations, object instances, and robot arms by adapting object-centric manipulation behaviors to new contexts through trajectory transformations based on known object poses, enabling the training for complex, long-horizon, and high-precision tasks.
@@
Additionally, DexGraspNet [261] featuring 1.32 million grasps of 5,355 objects by ShadowHand fills the blank of large-scale, diverse, and high-quality datasets for dexterous grasping. It includes over 200 diverse grasps per object, validated for physical stability in simulation, enabling more effective IL and benchmarking of robotic manipulation algorithms to achieve human-like dexterity and grasping capabilities.
@@ VI Challenges and Future Directions in Imitation Learning-Based Dexterous Manipulation
Imitation learning-based dexterous manipulation poses unique challenges due to the inherent complexities of both IL and dexterous control. Despite significant advancements over the past decade, several challenges hinder its human-level dexterity and real-world applicability. This section discusses these challenges from multiple perspectives and explores future research directions.
@@ VI-A Data Collection and Generation
Data collection and generation for imitation learning-based dexterous manipulation pose several challenges, including heterogeneous data fusion, data diversity, high-dimensional data sparsity, and data collection costs.